OpenStudy (anonymous):

1. Two friends leave school at the same time, Sarah is heading due north and Beth is heading due east. One hour later they are 10 miles apart. If Sarah had traveled 8 miles from the school, how many miles had Beth traveled? A. 8 miles B. 6 miles C. 10 miles D.4 miles

OpenStudy (anonymous):

bet you have the famous 3 - 4 - 5 right triangle here

OpenStudy (anonymous):



OpenStudy (anonymous):

12.8

OpenStudy (anonymous):

find the third side via pythagoras or just remember the mother of all right triangles 3-4-5

OpenStudy (anonymous):

oh no not 12.8

OpenStudy (anonymous):

you are looking for a leg, not the hypotenuse

OpenStudy (anonymous):

so subtract?

OpenStudy (anonymous):

\[8^2+?^2=10^2\] solve via \[?=\sqrt{10^2-8^2}\]

OpenStudy (anonymous):

or, remember a 3 - 4 - 5 right triangle multiply all by 2 get 6 - 8 - 10

OpenStudy (anonymous):

6

OpenStudy (anonymous):

what about this one?